A spacecraft is going from the earth to the moon. Find such a position from the earth over there The gravitational force is zero. Given
Earth mass = 6.0 × 10^24 kg. Moon mass = 74 × 10^22 kg; The distance between the center of the earth and the center of the moon = 3.8 x 10^8m
1
Expert's answer
2021-10-11T08:55:19-0400

Lets set the origin at the Earth. Then the force at distance "r" from the Earth is given as follows:


"F_1 = G\dfrac{M_Em}{r^2}"

where "G = 6.67\times 10^{-11}\dfrac{m^3}{kg\cdot s^2}" is the gravitational constant, "M_E" is the mass of the Earth, "m" is the mass of the ship.

The force from the Moon at the same point is given as follows:


"F_2 = G\dfrac{M_Mm}{(d-r)^2}"

where "d" is the distance between the center of the Earth and the center of the Moon. The total gravitational force is 0 if:


"F_1=F_2\\""G\dfrac{M_Em}{r^2} = G\dfrac{M_Mm}{(d-r)^2}\\
\dfrac{M_E}{r^2} = \dfrac{M_M}{(d-r)^2}"

Expressing "r", find:


"(M_E-M_M)r^2 - 2dM_Er + M_Ed^2 = 0"

Substituting the values and solving the quadratic equation, obtain:


"r\approx 2.8\times 10^8m"

Answer. "2.8\times 10^8m".